### v2

# For and some appropriate functions in [2] the authors considered the system of equations

1. Introduction
Fourth order parabolic PDE’s have remained an interesting topic of research for many decades. It is known that fourth order parabolic PDE’s are helpful in modelling various natural phenomenon and has found its application in many engineering problems such as, the free vibrations in beams or shafts [6], the growth of epitaxial thin film [12], thermal grooving by surface mechanism [18], and is also used for noise removal and edge preservation in image processing [25], to name a few. Recently research activities on inverse problems related to fourth order parabolic PDE has grown up substantially. Let us mention briefly a few recent works related to some inverse problems associated with fourth order parabolic equations.
For and some appropriate functions in [2] the authors considered the system of equations
and analysed the inverse problem of identifying the time-dependent coefficient called the Mullins’ coefficient, from the knowledge of the data where is any fixed given point and is the solution of the considered PDE. Subsequently, in [1], the author considered a more general fourth order parabolic PDE and investigated the inverse source identification problem associated with it. More precisely, the author considered the inverse problem of identification of the source function from the knowledge of final time observation or the time-integral observation for where is the solution of
for some given functions
As a follow up to the above developments, in a recent work [11], the authors have extended the investigation from integer-order derivative to fractional order setting. They considered an inverse problem of identifying time-dependent source term from the final value associated with a time-fractional fourth order parabolic PDE as follows:
where is a surface diffusion coefficient, is some appropriate function. Here (and throughout the paper) is the Caputo fractional derivative of order (cf. [20]) given by
where is the standard Gamma function, that is,
Note that forward problem associated with such time-fractional fourth order parabolic equations have been studied by many authors in various context, see for instance [7, 10].
It is important to recall that over the last few decades the research on time-fractional backward heat conduction problems (TFBHCP) have gained substantial attention due to its many practical applications. Since the work in [14], the literature on TFBHCP and its various variants have grown remarkably (see e.g., [9, 13, 23]). Considering the importance of fractional-order derivative, it is desirable to study various inverse problem associated with time-fractional fourth order parabolic equations. Recently some works on source identification problem associated with time-fractional fourth order parabolic equations have been reported (see e.g. [11]), as mentioned earlier. However, to the best of our knowledge, the backward problem for time-fractional fourth order parabolic equations has not been explored yet and thus, this paper intends to contribute in that direction.
In this paper we are interested in the inverse problem of retrieving the initial value from the knowledge of final value and the source function, associated with a time-fractional fourth order parabolic PDE. More precisely, we consider the following: Let (see Remark 3.5 for this restriction on ) be a bounded domain with sufficiently smooth boundary and be fixed. Let We consider the system
where and is the Caputo fractional derivative as defined in (1.1). The inverse problem that we consider is:
(IP): retrieve the initial value from the knowledge of and where is a solution of (1.2).
Similar to TFBHCP, the inverse problem to recover from the knowledge of associated with (1.2) is an ill-posed problem in the sense that a small perturbations in the data may lead to a large deviation in the sought initial value (see Example 3.9). Further, the problem of recovering for is stable with respect to the perturbations in the data (see Remark 3.8), possessing the properties similar to that of TFBHCP.
In practice, we shall have access to some measured data, and hence noise in the data is inevitable. We assume that for and are the noisy data satisfying
where is a fixed known constant that will be specified later (see Proposition 3.4).
Thus, the inverse problem at hand is to recover from the knowledge of and satisfying (1.3). Since the problem is ill-posed, in order to obtain stable approximations, we have to employ some regularization methods (see e.g., [4, 19]). We obtain stable approximations by the quasi-boundary value method (QBVM) (see e.g., [8, 23]), it’s modifications (MQBVM) (see e.g., [9, 15, 24]) and the Fourier truncation method (FTM) (see e.g., [13, 16]). Just to give some insight, in QBVM we perturb the final value in (1.2) by adding the initial value multiplied by the regularization parameter, that is, we consider the resulting perturbed term as where is the regularization parameter. For the MQBVM we consider the resulting perturbed term as where is the regularization parameter and . Note that in literature, the case has been studied extensively for various problems (see e.g., [24]), and is known by the name modified quasi-boundary value method. In FTM, we consider the truncated version of the Fourier expansion of the concerned function and the level at which we truncate plays the role of regularization parameter.
It is known in regularization theory that unless we assume some source condition (see e.g., [4, 19]) on the unknown, the convergence can be arbitrarily slow [22], that is, we can not obtain any convergence rate. In view of this, we assume that the unknown belongs to some source sets which are subsets of some Sobolev spaces where indicates the smoothness property, and then obtain error estimates and hence deduce the convergence rates of the approximations by the considered regularization methods. We shall provide both the apriori and the aposteriori parameter choice strategies and obtain the corresponding error estimates for all of them. As an important implication, we observe that for the FTM, the convergence rates in both the apriori and aposteriori cases is of the form for all where is the noise as considered in (1.3). However, for the QBVM and MQBVM, for both the apriori and aposteriori cases, the convergence rates possesses saturation effect. For the QBVM the best rates possible for the apriori and aposteriori cases is and for MQBVM the best rates possible for the apriori and aposteriori cases are and respectively, for Further, we observe that, if belongs to certain interval then the rates of FTM, QBVM, MQBVM are all the same for apriori case, and for certain values of the aposteriori rates for FTM and MQBVM are also the same. Moreover, we shall see that for the considered source sets the rates obtained by the FTM is always order optimal for all but the same is not the case for QBVM and MQBVM (see Remark 5.9).
The remainder of the paper is organised as follows: In Section 2 we list some notations, important facts, results about the Mittag-Leffler function and some basic result on calculus, that will be used throughout. In Section 3 we mainly discuss the relation between the sought initial value and the data, that is, the final value and the source function. We show by an example the ill-posedness of the considered inverse problem of retrieving the initial value. We also establish a conditional stability result. Section 4 is all about the QBVM and MQBVM. We provide both apriori and aposteriori parameter choice strategies and obtain the corresponding error estimates. Section 5 is about the FTM. We systematically define the regularized solutions and provide both apriori and aposteriori parameter choice strategies and obtain the corresponding convergence rates. Moreover, we discuss about the optimality of the obtained rate.
3. Preparatory results
In order to establish a relation between the sought initial value and the data, that is, the source function and the final value, we first consider the solution representation of a forward problem. We consider the PDE
where and
Let satisfies (3.1). Then following [11], we have
where recall that and
Definition 3.1.
Any satisfying (3.2) is said to be a solution of (3.1).
Clearly, given by (3.2) is indeed a solution of (3.1). Now suppose that it is given Then, we obtain
which implies
In order to reduce the notational complexity, for any and we define the following
where we recall that and
Thus, from (3.3), we have
and hence from (3.2), we have
Definition 3.2.
Let and Any that satisfies (3.6) is said to be a solution of the backward problem (1.2).
Before proceeding further, let us put in place some technical results.
Proposition 3.3.
Let Let and be as defined in (3.4). Then the following holds:
For all we have
For all we have
Proof.
(i). We have
(ii). The proof follows from (i).
Proposition 3.4.
Let Let and Let and where is as defined in (3.4). Then the following holds:
We have and
We have
Proof.
First we observe that
Now the proof follows.
We are now in a position to state and prove the necessary and sufficient condition for the existence and uniqueness of the solution of backward problem (1.2).
Theorem 3.6.
Let and The backward problem associated with (1.2) has a unique solution if and only if
and the solution is given by
Proof.
From Definition 3.2 and the solution representation (3.2) of the solution of (3.1), we observe that if as defined in (3.7), belongs to for all , then it is indeed a solution of the backward problem (1.2). Thus, it remains to ensure that for
Towards this, for by invoking Theorem 2.1, we first observe that
which is independent of Hence by Propositions 3.3 (i) and 3.4 (i), it follows that
For observe that and Thus, if and only if
Theorem 3.7.
For let and be the corresponding solutions for the backward problem associated with (1.2). Then for there exists a constant depending on such that
Proof.
From (3.7), we have
where recall that is as in Proposition 3.4. Now the proof follows from Propositions 3.3 and 3.4.
The next question one would like to ask is, if is a solution of the backward problem (1.2), then whether is stable under perturbations in and The answer is in the negative as is seen in the following example.
Example 3.9.
(Ill-posed)
For and let and Corresponding to the source function and the final value let denote the initial value associated with the backward problem (1.2), and for the source function and the final value let denote the initial value associated with the backward problem (1.2). Then from the representation (3.7), it follows that and Now we observe that as but
as
We now obtain an equivalent formulation of the inverse problem of retrieving the initial value from the final value and the source function.
Let , and be as in Proposition 3.4. Let be the solution of the backward problem (1.2). Then from (3.5) and Proposition 3.4 (i), we observe that the inverse problem of recovering the initial value from the knowledge of and is same as solving the linear operator equation
where is defined by
From Theorem 2.1 it follows that and thus it follows that is a compact linear operator of infinite rank. Therefore, solving (3.9) is an ill-posed problem. Moreover, if
then (3.9)
has a unique solution given by
where is as defined in (3.5) and is as defined in (3.4).
We now recall definitions of some Sobolev spaces that are relevant in our analysis.
For let
It is known that is a Hilbert space with respect to the inner-product
and the corresponding norm is given by
Note that and
For some we consider the following set
These sets will be the source sets that are of interest in our analysis.
We end this section by proving a conditional stability estimate.
Theorem 3.10.
Let and be as in Proposition 3.4. Let be the unique solution of (3.9). If for some then there exists a constant depending on and such that
Proof.
Since is a solution of (3.9), we have , where are as in (3.5), that is,
Thus,
Now, from Theorem 2.1 we have
where Thus,
using the fact that
we have
that is,
Now the proof follows.
4. Quasi-boundary value method and its modification
We have seen in the preceding section that the considered inverse problem of retrieving the initial value from the knowledge of final value and the source term, is an ill-posed problem in the sense that a small perturbations in the data and may lead to a large deviation in the sought solution. But, in practice, we shall have knowledge of the data that are corrupted by some noise. Thus, we have to employ some regularization method to obtain stable approximations. In this section we shall employ the quasi-boundary value method and its modifications, and obtain error estimates for apriori as well as aposteriori parameter choice strategies.
For and we consider the following equation
where and Note that will play the role of regularization parameter.
By separation of variable method, we know that the solution of (4.1) is of the form
where ’s are constants to be determined. Now using the condition we obtain
where is as defined in (3.4). Therefore, the solution of (4.1) is given by
Now taking the noisy data and in place of the exact data and , respectively, the solution of (4.1) is given by
where is as defined in (3.4).
LEMMA 4.1.
Let and . For , let and The following holds:
There exists a constant depending on such that
For we have
where is as in Proposition 3.4.
Proof.
(i). Let Since it follows that for all
Let From Theorem 2.1, we have
where Thus, we have
Now the result follows from Lemma 2.3 (i).
(ii).
The proof follows from Proposition 3.4 and Lemma 4.1 (i).
Thus, it follows that for every and the equation (4.1) has a unique solution given by (4.2). In particular, we have
Note that these are the candidates for the regularized solutions in this section.
For and from (4.3), we have
Theorem 4.3.
Let be as in Proposition 3.4, be as in Lemma 4.1 and . For let and be such that
For let and be as defined in (4.4) and (4.5), respectively. Then there exists a constant depending on such that
Proof.
We have
Now the proof follows from Lemma 4.1 (i) for , and by taking for
For recall the definition of the source set as given in (3.11).
Theorem 4.5.
Let and for let be as in (4.4), be as in Lemma 4.1 and be the unique solution of (3.9). If for some then there exist constants depending on and depending on such that
Proof.
Recall that where is as in (3.5). Thus, we have
where
Now from Theorem 2.1, we observe that
where
Case 1. Let
Then by Lemma 2.3 (ii), it follows that
where
Case 2. Let
Then
where
Now we are in a position to state one of the main result of this section whose proof follows by combining the estimates obtained in Theorems 4.3 and 4.5, respectively. This result provides the rates of convergence for the approximations by regularized solutions obtained by modified quasi-boundary value method for all under an apriori parameter choice strategy.
Theorem 4.6.
For and let be as in Theorem 4.3. Let be as in Theorem 4.5. Let be as in Theorems 4.3 and 4.5, respectively. Then we have the following:
For and the choice there exists a constant depending on such that
For and the choice there exists a constant depending on such that
For and the choice there exists a constant depending on , such that
For and the choice there exists a constant depending on , such that
In Theorem 4.6 we have obtained the rates by choosing the regularization parameter in a certain manner that requires the apriori knowledge of This means that one has to apriori know the smoothness of the unknown But that is not always feasible and is too restrictive to assume. In order to overcome this, we provide an aposteriori parameter choice strategy in the subsequent analysis.
4.1. Error estimates with aposteriori parameter choice
For and let be as in Proposition 3.4, that is, where is as defined in (3.4). Also, recall that the considered inverse problem is formulated as solving an operator equation where is as defined in (3.10).
We now state a result which has been used frequently in the literature. Since the context in which we are using is new, we include only the key steps in the proof for the sake of completeness.
LEMMA 4.7.
For and let be as in (4.5) and be as defined above with Let
Then the following holds:
is continuous and increasing.
and
Proof.
(i). The continuity follows easily by observing that
In order to show that is increasing, we first observe that for the function is increasing for Since there exists some such that Using these facts, it follows that is indeed increasing.
(ii). From the expression of it follows that Next, from Theorem 2.1, we obtain
where This implies that
However, we also observe that
Now the proof follows by taking the limit as
LEMMA 4.8.
Let be the solution of (3.9) and suppose that for some For let be as defined in Lemma 4.7. Let be given and be fixed. Then for every small enough, with there exist unique and such that
and the following result holds:
For and there exists a constant depending on such that
For and there exists a constant depending on such that
For we have
Proof.
From Proposition 3.4 (ii), we first note that for
Let Now the existence of unique satisfying and follows from Lemma 4.7 by observing that Thus, we have
Thus,
where
Now, from Theorem 2.1, it can be verified that
where and
(i). Let and
Then we have
Thus, in this case,
where Therefore,
Thus, we have
(ii). Let and
In this case, we have
where Thus,
and hence
(iii). Let Recall that there exists a unique such that Thus, we have
Therefore,
and hence
Therefore,
Thus, from Theorem 4.3 and Lemma 4.8, we have the following result.
Theorem 4.9.
Let and be as in Lemma 4.8. Let and be as in (4.5) and (4.4), respectively. Let be the solution of (3.9) and be as in Theorem 4.3 and Lemma 4.8. If for some then the following result holds:
For there exists a constant depending on such that
For there exists a constant depending on such that
For there exists a constant depending on such that
We now obtain estimate for and
Theorem 4.10.
Let be as in Theorem 4.9, be as in (4.4) and be as in Theorem 3.10. Let be the solution of (3.9). If for some then there exists a constant depending on such that the following holds:
For
For
Proof.
From the definition of (see (3.10)), we have
and hence
(i). Let
In this case, we have
Now, recall that
Thus,
This shows that and Thus, from Theorem 3.10, we have
where
(ii). Let
Now repeating the arguments of (i) and using the fact , we obtain
Also, it follows that and Now, the proof follows by invoking the estimate in Theorem 3.10.
Now we are in a position to state the main result of this section, whose proof follows by combining the results in Theorems 4.9 and 4.10, respectively.
Theorem 4.11.
Let and be as in Lemma 4.8. Let be the solution of (3.9). If for some then the following holds:
For we have
For we have
For we have
5. Fourier truncation method
In this section we obtain regularized approximations by truncating the Fourier expansion of appropriate functions. The level at which the series expansion is truncated plays the role of regularization parameter. We shall provide both the apriori and aposteriori parameter choice strategies and derive the corresponding convergence rates of the approximation. We shall see that the rates obtained by this method are similar for both the apriori and aposteriori parameter choice. Moreover, we shall see that the rates obtained by this method does not possess the saturation effect which is in contrast to the rates obtained by quasi-boundary value method and its modifications, as obtained in the previous section.
Throughout this section we assume the following: Let be as in Theorem 4.3, that is, , and
where
Let , be as defined in (3.4) and be the solution of (3.9). Recall that solving (3.9) is an ill-posed problem with respect to the perturbations in the data and . Thus, we aim to obtain stable approximations for the sought initial value by the Fourier truncation method. Towards this, for we define
Here the natural number plays the role of regularization parameter. We shall choose in terms of so that under certain source condition we can derive the error estimates in terms of the noise
First, we obtain an estimate for For recall the source set as defined in (3.11).
Theorem 5.1.
Let be the solution of (3.9) and for let be as defined in (5.2). If for some then we have
Proof.
Since is a solution of (3.9), recall from (3.5) that, we have where Thus, from the definition of we have
that is,
We now obtain bound for
Theorem 5.2.
For and let and be as defined in (5.2) with satisfying the noise level as considered in (5.1). Then there exists a constant depending on such that
Proof.
First we observe that from Theorem 2.1, we have
where Now from the definition of and we have
where Therefore,
We are now in a position to state and prove one of the main result of this section, that is, the apriori parameter choice strategy and the corresponding error estimates. From now onwards we shall use the notation to denote the greatest integer not exceeding where
Theorem 5.4.
Let and be as in (2.2). For and let be as in Theorem 5.2. For let be as in Theorem 5.1. Let be as in Theorem 5.2. Then for the choice
there exists a constant depending on such that
Proof.
Combining the estimates obtained in Theorems 5.1 and 5.2, we have
For let
Then it is easy to verify that attains its minimum at satisfying
That is, we have,
We choose the regularization parameter as that is,
Then with this choice of we have,
where
This completes the proof.
In Theorem 5.4 we have provided a parameter choice strategy and obtained the corresponding convergence rate. However, we observe that the choice of the regularization parameter depends on which is related to the smoothness of the unknown Therefore, the parameter choice strategy will work only if we know apriori the smoothness of But, in practice, it is almost always impossible to know the smoothness apriori. Thus, to overcome this situation, we provide the aposteriori parameter choice that only requires knowledge of the noise and the observed data
5.1. Error estimates with aposteriori parameter choice
Recall that from Proposition 3.4, it follows that for and we have that where is as defined in (3.4). For we let
We define
Then it is easily seen that as
Let Then, for sufficiently small, there exists such that
Let be the first natural number that satisfies (5.3). This will be the regularization parameter that we are looking for. Note that, clearly the choice of such depends only on the known quantities,
We now obtain a bound for in terms of
LEMMA 5.5.
Let be as in (2.2) and be the exact data. For let the noisy data and be such that it satisfies (5.1). Let be fixed, be the smallest natural number satisfying (5.3) and be as in Theorem 2.1. Let be the solution of (3.9). If for some then there exists a constant depending on such that
Proof.
First we note that
Thus,
where Thus,
Now, observe that
Therefore, we have
Thus,
and hence
We now obtain estimate for
Theorem 5.6.
Let be the solution of (3.9) and for some Let be as in Lemma 5.5 and be as defined in (5.2). Let be as in Theorem 2.1. Then there exists a constant depending on such that
Proof.
Since is a solution of (3.9), recall that where is as given in (3.5), that is,
Thus, from the definition of we have
where
Now, recall that
and
Thus,
Thus, we obtain
that is,
where
We are now in a position to state and prove the main result of this section, in which we obtain the rates of convergence corresponding to the aposteriori choice of the parameter .
Theorem 5.7.
Let and be as in Lemma 5.5. Let be as in (5.2) and be as in (2.2). Let be the solution of (3.9). If for some then there exists a constant depending on such that
Proof.
From the estimate obtained in Theorem 5.2, we have
where and is as in (2.2).
Thus, from Theorem 5.6, we have
where